Two bicyclists, 48 miles apart, begin riding toward each other on a long straight avenue. One cyclist travels 15 miles per hour and the other 25 miles per hour. At the same time, Spot (a greyhound), starting at one cyclist, runs back and forth between the two cyclists as they approach each other. If Spot runs 35 miles per hour and turns around instantly at each cyclist, how far has he run when the cyclists meet?

Answer by richwmiller(17219)

The bike from b had a(n) 0 hour headstart
The distance traveled in those 0 hours is 0 miles
The distance to go 48 miles
d=r*t+R*t
48=15*t+25*t
48=40*t
t=1.2 hours
The total time for bike from b to meet the bike from a is 1.2 hours
d=R*t
d=25*1.2
d=30.0+0
d=30.0 miles from b at 25 mph
d=r*t
d=15*1.2
d=18.0 miles from a at 15 mph
For out purpose it doesn't matter which direction Spot is running.
The cyclist are moving for 1.2 hours and so is Spot.
1.2*35=42 Spot has run 42 miles
I don't think greyhounds have that stamina!